OpenStudy (anonymous):

A ladder leans against a house at a 60° angle to the ground.

OpenStudy (anonymous):

If the ladder extends to a length of 132 inches, what is the height of the house rounded to the nearest hundredth of an inch? Type the numeric answer only in the box below. Numerical Answers Expected!

OpenStudy (rane):

wht is your ans?

OpenStudy (anonymous):

You will use here: \[\sin(\theta) = \frac{Perpendicular}{Hypotenuse}\]

OpenStudy (anonymous):

So: \[\large \color{green}{\sin(60^{\circ}) = \frac{x}{132}}\] Solve this for x..

OpenStudy (rane):

so u times them

OpenStudy (anonymous):

I had got it wrong. My answer was 114.31

OpenStudy (rane):

wht is the correct or given ans?

OpenStudy (anonymous):

Just multiply value of sin(60) with 132, you will have x..

OpenStudy (rane):

thats exactly my answer too @marylou004

OpenStudy (anonymous):

Yep.. Right..

OpenStudy (rane):

its 132sin60

OpenStudy (anonymous):

No, sin(60) with 132..

OpenStudy (anonymous):

You are getting it right or not?

OpenStudy (anonymous):

I am confused what is the numerical answer? The answer I had of 114.31 was wrong.

OpenStudy (anonymous):

Is it wrong?? Have you tried for 114.84??

OpenStudy (rane):

then wht is the correct ans?

OpenStudy (anonymous):

No only tried 114.31

OpenStudy (anonymous):

114.32=x=EG=Height of building EF=66

OpenStudy (anonymous):

From other way I got 114.84, but not sure it is right or not.. 114.31 should be right..

OpenStudy (anonymous):

Ok

OpenStudy (rane):

u shouldn't be using cosine rule because the hypoteneius is opposite to the angle not adjacent

OpenStudy (rane):

@pasta

OpenStudy (rane):

so, the ans should be 114.31

OpenStudy (anonymous):

So did we decide the correct answer? I know it isn't 114.31 because I got it wrong.

OpenStudy (anonymous):

Once try for 114.84..

OpenStudy (anonymous):

ok

OpenStudy (rane):

thats really weird then, then how abt try 66, even though i doubt it bt just give it a go

OpenStudy (anonymous):

@RANE your right 114.32

OpenStudy (rane):

ok @pasta bt apparently its wrong as @marylou004 said

OpenStudy (anonymous):

I sais 114.31 was wrong.

OpenStudy (rane):

yeah i know @pasta just rounded off but its still the same

OpenStudy (rane):

did u try 66

OpenStudy (anonymous):

114.32 is correct. I guess I just forgot to round it. Thanks to everyone.

OpenStudy (rane):

okk....i got really confused

OpenStudy (anonymous):

Welcome..

OpenStudy (rane):

my pleasure

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

:)